# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v6

@@ VI-D Extensions
As a sub-field of LLMs and natural language understanding research, many studies in these fields have been adopted for text-to-SQL tasks, advancing its development [149, 126]. However, text-to-SQL research can also be extended to the larger scope studies of these fields at meanwhile. For instance, SQL generation is a part of code generation. The well-designed approaches in code generation also obtain promising performance in text-to-SQL [79, 80], performing generalization across various programming languages. The potential extension of some tailored text-to-SQL frameworks to NL-to-code studies can also be discussed. For instance, frameworks integrating execution output in NL-to-code can also achieve solid performance in SQL generation [7].
@@
An attempt to extend execution-aware approaches in text-to-SQL with other advancing modules [70, 28] for code generation is worth discussing. From another perspective, we previously discussed that text-to-SQL can enhance LLM-based question-answering (QA) by providing factual information. The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [150, 151]). Construct the factual knowledge with database structure, and then incorporate the text-to-SQL system to achieve information retrieval, which can potentially assist further QA with more accurate factual knowledge [152]. More extensions of text-to-SQL studies are expected in future work.

### v7

@@ VI-D Extensions
As a sub-field of LLMs and natural language understanding research, many studies in these fields have been adopted for text-to-SQL tasks, advancing its development [wei2022chain, wang2023selfconsistency]. However, text-to-SQL research can also be extended to the larger scope studies of these fields at meanwhile. For instance, SQL generation is a part of code generation. The well-designed approaches in code generation also obtain promising performance in text-to-SQL [ni2023lever, chen2024selfdebugging], performing generalization across various programming languages. The potential extension of some tailored text-to-SQL frameworks to NL-to-code studies can also be discussed. For instance, frameworks integrating execution output in NL-to-code can also achieve solid performance in SQL generation [pourreza2023dinsql].
@@
An attempt to extend execution-aware approaches in text-to-SQL with other advancing modules [hong2024knowledgetosql, dong2023c3] for code generation is worth discussing. From another perspective, we previously discussed that text-to-SQL can enhance LLM-based question-answering (QA) by providing factual information. The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [luo2023chatkbqa, li2024flexkbqa]). Construct the factual knowledge with database structure, and then incorporate the text-to-SQL system to achieve information retrieval, which can potentially assist further QA with more accurate factual knowledge [xiong2024interactive]. More extensions of text-to-SQL studies are expected in future work.
[Uncaptioned image] Zijin Hong is a first-year Ph.D. student in the DEEP Lab, Department of Computing, The Hong Kong Polytechnic University, under the supervision of Prof. Xiao Huang. He holds a dual bachelor’s degree in Applied Mathematics with Mathematics from Jinan University and the University of Birmingham. His research interests include machine learning, large language models, and their applications, particularly LLM-based text-to-SQL. He has published papers and served as a reviewer for venues such as ACL, EMNLP, NAACL, AAAI, ICLR, and IEEE TKDE.
[Uncaptioned image] Zheng Yuan is a second-year Ph.D. student at the Department of Computing, The Hong Kong Polytechnic University. He received a bachelor degree in Computer Science and Technology from Shenzhen University in 2021. He is currently a group member in the DEEP Lab supervised by Prof. Xiao Huang at the Hong Kong Polytechnic University. His research interests focus on large language models and their applications in data-driven scenarios, including retrieval-augmented generation, text-to-SQL, and recommender systems.
[Uncaptioned image] Qinggang Zhang is currently a Postdoctoral Fellow at the Deep Lab, Department of Computing, The Hong Kong Polytechnic University, Hong Kong SAR. He received his Ph.D. and B.Eng. degrees from The Hong Kong Polytechnic University and Northwestern Polytechnical University, respectively. His research interests include KGs, LLMs, RAG, and Text-to-SQL. He has published over 20 papers while serving as a reviewer for NeurIPS, ICML, ICLR, KDD, IEEE TKDE, and IEEE TPAMI.
[Uncaptioned image] Hao Chen is an assistant professor at the Faculty of Data Science, City University of Macau. Before that, he worked as a postdoctoral research fellow at The Hong Kong Polytechnic University. He received his Ph.D degree from Beihang University, China in 2022. He has published over 40 papers, such as TKDE, NeurIPS, ACL, SIGKDD, WWW, and SIGIR, etc. He received the Best Paper Award Honorable Mention at SIGIR 2023. His research interests include recommender systems, graph neural networks, and large language models.
[Uncaptioned image] Junnan Dong is the final-year Ph.D. candidate in the department of computing, Hong Kong Polytechnic University. He focuses on Large Language Models, Knowledge Graphs and Graph Retrieval Augmented Generation for advanced and knowledgeable domain-specific reasoning. He has published more than ten papers on top-tier conferences, including NeurIPS, ACL, WWW, KDD, SIGIR and WSDM. He has been continuously serving as a program committee member for NeurIPS, ICML, ICLR, ACL and WWW.
[Uncaptioned image] Feiran Huang received the Ph.D. degree from the School of Computer Science and Engineering, Beihang University, Beijing, China, in 2018. He is currently a professor with the College of Information Science and Technology, Jinan University. He has over 100 publications appearing in top conferences and journals, such as SIGIR, KDD, WWW, MM, TIP, IJCV, and TKDE. He received 5 best paper awards, such as the Best Paper Award Honourable Mention in SIGIR 2023 and Best Paper Award Runner Up in PAKDD 2023. He holds over 50 US, Chinese, and international granted patents. He is an editorial board member of IEEE Transactions on Affective Computing, ACM Transactions on Recommender Systems, etc. His research interests include recommender systems, social networks, sentiment analysis, and large language models.
[Uncaptioned image] Xiao Huang is an Assistant Professor in the Department of Computing at The Hong Kong Polytechnic University. He earned his Ph.D. in Computer Engineering from Texas A&M University in 2020, an M.S. in Electrical Engineering from the Illinois Institute of Technology in 2015, and a B.S. in Engineering from Shanghai Jiao Tong University in 2012. His scholarly contributions are highly regarded within the academic community, amassing over 4,500 citations. He received the Best Paper Award Honorable Mention at SIGIR 2023. He has successfully led or completed seven research projects as Principal Investigator.